Ariana took 2 hours longer to drive 360 miles on the first day of a trip than she took to drive 270 moles on the second day. If her speed was the same on both days what was the driving time each day?
Answer by checkley71(8403)

360-270=90 MILES DIFFERENCE
90/2=45 MPH. IS THE SPEED
PROOF
360/45=270/45+2
8=6+2
8=8